# REVEAL-IT: REinforcement learning with Visibility of Evolving Agent poLicy for InTerpretability

### v1

@@ 2 Related Works
Explainability for RL. Explainable RL methods can be broadly categorized into two groups: post hoc and intrinsic approaches (Puiutta & Veith, 2020; Heuillet et al., 2020). Post hoc explanations are provided after the model execution, whereas intrinsic approaches inherently possess transparency. Post hoc explanations, such as the saliency map approach (Greydanus et al., 2017; Mott et al., 2019), often rely on correlations. However, as we mentioned earlier, conclusions drawn based on correlations may be unreliable and fail to answer causal questions. On the other hand, intrinsic explanations can be achieved using interpretable algorithms, such as linear regression or decision trees, but the limited modeling capacity of these algorithms may prove insufficient in explaining complex behaviors (Puiutta & Veith, 2020).
Explanation in real world. It is worth noting that the practical implementation of causal RL in real-world applications remains limited. To make it more widely applicable, we must carefully examine reality’s challenges.  Dulac-Arnold et al. (2020); Paduraru et al. (2021) identified nine critical challenges that are holding back the use of RL in the real world: limited samples; unknown and large delays; high-dimensional input; safety constraints; partial observability; multi-objective or unspecified reward functions; low latencies; offline learning; and the need for explainability. In this paper, we will focus on how to address the explanation of RL in a high-dimensional environment.
@@ 3 Preliminaries
Our framework comprises of two parts. One is the RL policy visualization module that shows the policy update in training and policy activation in the test. This module is based on fully connected neural networks’ well-established node-link diagram representation. The other is a GNN-based explainer, trained to analyze the visualized results in the first module and underly the important update in the training phase. We can further explore the relationship between different training tasks and the contribution of a specific training task/phase to the final performance. Now, we provide more details on these two parts:
@@ 3.1 Reinforcement Learning and Visualization
Reinforcement Learning. The problem of controlling an agent can be modeled as a Markov decision process (MDP), denoted by the tuple$\{\mathcal{S},\mathcal{A},P,R,\mu_{0},\gamma\}$, where $\mathcal{S}$ and $\mathcal{A}$ represent the set of state space and action space respectively, $P:\mathcal{S}\times\mathcal{A}\times S\rightarrow[0,1]$ is the transition probability function that yields the probability of transitioning into the next state $s_{t+1}$ after taking an action $a_{t}$ at state $s_{t}$, and $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function that assigns the immediate reward for taking an action $a_{t}$ at state $s_{t}$. $\mu_{0}:\mathcal{S}\rightarrow[0,1]$ is the probability distribution that specifies the generation of the initial state, and $\gamma\in[0,1]$ denotes the discount factor. We can use a deep RL algorithm to train the agent (e.g., DQN (Mnih et al., 2015), A2C (Mnih et al., 2016), SAC (Haarnoja et al., 2018b), PPO Schulman et al. (2017), etc.). In the experiments of this paper, we use DQN and PPO as the fundational RL methods. Note that any online RL algorithm can be accepted.
@@
4:         Randomly sample training task sequence $Seq_{t}$ in $\mathcal{D}_{T}$
@@
6:         Sample training tasks $Seq_{t}$ in terms of $y_{T}$
@@
8:     RL exploration and collect data to $\mathcal{B}$ following $Seq_{t}$triangleright RL exploration
9:     Update $\pi_{\theta}$ using samples from $\mathcal{B}$
10:     Visualize policy update and save update information into $\mathcal{D}$triangleright Visualization
11:     Train GNN to predict $y_{T}$ based on $\mathcal{D}$triangleright Train GNN explainer
@@
3.2 Background on GNN-based Explainer
The $l$-th layer of GNN explainer performs three essential computations: MESSAGE, AGGREGATE and UPDATE. The model first calculates the messages between every pair of nodes by $m_{ij}^{l}=MESSAGE(h_{i}^{l-1},h_{j}^{l-1},r_{ij}$, where $h_{i}^{l-1}$ and $h_{j}^{l-1}$ denote the node representations of $v_{i}$ and $v_{j}$ obtained from the previous layer $l-1$, respectively, and $r_{ij}$ represents the relation between them. Next GNN aggregates the messages for each node $v_{i}$ by $m_{i}^{l}=AGGREGATE(m_{ij}^{l}|v_{j}\in\mathcal{N}(v_{i})),$ where $\mathcal{N}(v_{i})$ refers to the neighborhood of $v_{i}$. Finally, the GNN explainer updates the node representation from the previous layer by $h_{i}^{l}=UPDATE(m_{i}^{l},h_{i}^{l-1})$. The final embedding for node $v_{i}$ after $L$ layers of computation is $z_{i}=h_{i}^{L}$, which is then used for node classification, i.e., the understanding the critical nodes in the RL agent’s evaluation is a crucial pre-requisite for determining the significance of weights updating.
4 REVEAL-IT
@@
where $G_{X}$ represents the explanatory subgraph reflecting the important edges (corresponding to policy updates), and $\Delta G$ is the remaining graph containing unimportant edges. When explaining a graph of visualized policy updates, we denote the optimal subgraph by $G_{X}^{m}$. When $G_{X}^{m}$ is used as input to $\Phi$, it can retain the original prediction. $G_{X}^{m}$ is optimal, as removing the features from it would result in a different prediction.
@@ 6 Conclusion and Future Work
Our research centers on the explanation of an RL agent’s training tasks and performance in a continuous control environment. This study aims to solve the explanation problem for RL in complex environments and tasks. enhancing transparency and interpretability in RL. This framework can promote trust and understanding of AI decision-making processes by explaining an RL agent’s training tasks and performance in complex environments. This increased transparency can lead to more responsible and ethical deployment of RL systems in various real-world applications, such as autonomous vehicles, healthcare diagnostics, and financial trading. Technological progressions possess the capability to augment industrial procedures, ameliorate efficacy, and mitigate hazards to human laborers across diverse domains.
Acknowledgements
This work is supported by the Air Force Office of Scientific Research under award number FA2386-23-1-4066. Case Number: 2024-0624 (original case number(s): AFRL-2024-2944).

### v4

@@ 2 Related Works
Explainability for RL.  Puiutta & Veith (2020); Heuillet et al. (2020) categorize the explainable RL methods into two groups. The first group is “Post hoc”, which aims to provide explanations after the model execution, while the other group is “intrinsic approaches”, which inherently possess transparency. Many post hoc explanations, like the saliency map approach (Greydanus et al., 2017; Mott et al., 2019), are based on correlations. This means that conclusions drawn from correlations may not be accurate and do not answer questions about what caused what. You can use interpretable algorithms, like linear regression or decision trees, to develop intrinsic explanations. However, these algorithms may not be able to model complex behaviors (Puiutta & Veith, 2020) fully. Our method proposes a new mechanism in the problem of explainability for RL from the perspective of the agent’s learning process, i.e., it can help to understand how the agent is learning in a given task or environment.
Explanation in real world.  Deng et al. (2023) provides a general discussion on the challenges posed by reality in real-world scenarios. The non-stationary and constant evolution in the real world requires the RL algorithm to be more robust to handle these changes.  Dulac-Arnold et al. (2020); Paduraru et al. (2021) found nine major problems with RL that make it hard to use in real life. These problems are small sample sizes, unknown or long delays, complex inputs, safety concerns, limited visibility, unclear or multiple reward functions, short delays, learning while offline, and the need to be able to explain. In this paper, we will address the limitations of generating explanations for RL in a high-dimensional environment from the perspective of the training tasks and the learning progress.
@@ 3 Preliminaries
Our framework comprises of two parts. One is the RL policy visualization module that shows the policy update in training and policy activation in the test. This module is based on fully connected neural networks’ well-established node-link diagram representation. The other is a GNN-based explainer, trained to analyze the visualized results in the first module and underly the critical update in the training phase. We can further explore the relationship between different training tasks and the contribution of a specific training task/phase to the final performance. Now, we provide more details on these two parts:
@@ 3.1 Reinforcement Learning and Visualization
Reinforcement Learning. The problem of controlling an agent can be modeled as a Markov decision process (MDP), denoted by the tuple$\{\mathcal{S},\mathcal{A},P,R,\mu_{0},\gamma\}$, where $\mathcal{S}$ and $\mathcal{A}$ represent the set of state space and action space respectively, $P:\mathcal{S}\times\mathcal{A}\times S\rightarrow[0,1]$ is the transition probability function that yields the probability of transitioning into the next state $s_{t+1}$ after taking an action $a_{t}$ at state $s_{t}$, and $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function that assigns the immediate reward for taking an action $a_{t}$ at state $s_{t}$. $\mu_{0}:\mathcal{S}\rightarrow[0,1]$ is the probability distribution that specifies the generation of the initial state, and $\gamma\in[0,1]$ denotes the discount factor. We can use a deep RL algorithm to train the agent (e.g., DQN (Mnih et al., 2015), A2C (Mnih et al., 2016), SAC (Haarnoja et al., 2018b), PPO (Schulman et al., 2017), etc.). In the experiments of this paper, we use DQN and PPO as the fundational RL methods. Note that any online RL algorithm can be accepted.
@@
3.2 Background on GNN-based Explainer
The $l$-th layer of GNN explainer performs three essential computations: MESSAGE, AGGREGATE and UPDATE. The model first calculates the messages between every pair of nodes by $m_{ij}^{l}=MESSAGE(h_{i}^{l-1},h_{j}^{l-1},r_{ij})$, where $h_{i}^{l-1}$ and $h_{j}^{l-1}$ denote the node representations of $v_{i}$ and $v_{j}$ obtained from the previous layer $l-1$, respectively, and $r_{ij}$ represents the relation between them. Next GNN aggregates the messages for each node $v_{i}$ by $m_{i}^{l}=AGGREGATE(m_{ij}^{l}|v_{j}\in\mathcal{N}(v_{i})),$ where $\mathcal{N}(v_{i})$ refers to the neighborhood of $v_{i}$. Finally, the GNN explainer updates the node representation from the previous layer by $h_{i}^{l}=UPDATE(m_{i}^{l},h_{i}^{l-1})$. The final embedding for node $v_{i}$ after $L$ layers of computation is $z_{i}=h_{i}^{L}$, which is then used for node classification, i.e., the understanding the critical nodes in the RL agent’s evaluation is a crucial pre-requisite for determining the significance of weights updating.
4 REVEAL-IT
@@
4:         Randomly sample training task sequence $Seq_{t}$ in $\mathcal{D}_{T}$;
@@
6:         Sample training tasks $Seq_{t}$ in terms of $y_{T}$;
@@
8:     RL exploration and collect data to $\mathcal{B}$ following $Seq_{t}$; $\triangleright$ RL exploration
9:     Update $\pi_{\theta}$ using samples from $\mathcal{B}$;
10:     Visualize policy update and save update information into $\mathcal{D}$$\triangleright$ Visualization
11:     Train GNN to predict $y_{T}$ based on $\mathcal{D}$; $\triangleright$ Train GNN explainer
@@
where $G_{X}$ represents the explanatory subgraph reflecting the important edges (corresponding to policy updates), and $\Delta G$ is the remaining graph containing unimportant edges. When explaining a graph of visualized policy updates, we denote the optimal subgraph by $G_{X}^{m}$. When $G_{X}^{m}$ is used as input to $\Phi$, it can retain the original prediction. $G_{X}^{m}$ is optimal, as removing the features from it would result in a different prediction. We will apply GNN explainer to
@@ 6 Conclusion and Future Work
Our research conducted centers on the explanation of an RL agent’s training tasks and performance in a continuous control environment. This study aims to solve the explanation problem for RL in complex environments and tasks. enhancing transparency and interpretability in RL. This framework can promote trust and understanding of AI decision-making processes by explaining an RL agent’s training tasks and performance in complex environments. This increased transparency can lead to more responsible and ethical deployment of RL systems in various real-world applications, such as autonomous vehicles, healthcare diagnostics, and financial trading. Technological progressions possess the capability to augment industrial procedures, ameliorate efficacy, and mitigate hazards to human laborers across diverse domains.